(x+2y-1)dx - (x+2y+1) dy =0
1
Expert's answer
2020-07-23T17:56:10-0400

"(x+2y-1)dx=(x+2y+1)dy"

 


ODE classification:

first-order nonlinear ordinary differential equation.

Step-by-step solution:

1-step:

Solve "x+2y(x)-1=(x+2y(x)+1)\tfrac{dy(x)}{dx}"

2-step:

Let "v(x)=x+2y(x)," which gives "\tfrac{dv(x)}{dx}=2\tfrac{dy(x)}{dx}+1:"


"v(x)-1=\tfrac{1}{2}(\tfrac{dv(x)}{dx}-1)(v(x)+1)"

3-step:

Solve for "\tfrac{dv(x)}{dx}:"


"\tfrac{dv(x)}{dx}=\tfrac{3v(x)-1}{v(x)+1}"

4-step:

Divide both sides by "\tfrac{3v(x)-1}{v(x)+1}" :


"\tfrac{\tfrac{dv(x)}{dx}(v(x)+1)}{3v(x)-1}=1"


5-step:

Integrate both sides with respect to x:


"\int" "\tfrac{\tfrac{dv(x)}{dx}(v(x)+1)}{3v(x)-1}dx=\int1dx"


6-step:

Evaluate the integrals:


"\tfrac{1}{9}(4log(3v(x)-1)+3v(x)-1)=x+C_{1}"


where "C_{1}" is an arbitrary constant.


Note:"W" is the product log function.

7-step:

Solve for "v(x):"


"v(x) = \tfrac{1}{3}(4W(- \tfrac{1}{4} \sqrt[4]{e^{9(x+C_{1})}})+1" )

or

"v(x) = \tfrac{1}{3}(4W(- \tfrac{1}{4}i \sqrt[4]{e^{9(x+C_{1})}})+1)"

or

"v(x) = \tfrac{1}{3}(4W( \tfrac{1}{4}i \sqrt[4]{e^{9(x+C_{1})}})+1)"

or

"v(x) = \tfrac{1}{3}(4W( \tfrac{1}{4} \sqrt[4]{e^{9(x+C_{1})}})+1)"


8-step:

Substitute back for "y(x) =" "\tfrac{1}{2}(-x+v(x)):"


Answer:

"y(x) = \tfrac{1}{6}(-3x+4W(- \tfrac{1}{4} \sqrt[4]{e^{9(x+C_{1})}})+1)"

or

"y(x) = \tfrac{1}{6}(-3x+4W(- \tfrac{1}{4}i \sqrt[4]{e^{9(x+C_{1})}})+1)"

or

"y(x) = \tfrac{1}{6}(-3x+4W( \tfrac{1}{4}i \sqrt[4]{e^{9(x+C_{1})}})+1)"

or

"y(x) = \tfrac{1}{6}(-3x+4W(\tfrac{1}{4} \sqrt[4]{e^{9(x+C_{1})}})+1)"



Differential equation solution:

"y(x)=\tfrac{2}{3}(W(-e^{\tfrac{9x}{4+C_{1}-1}})+1)+\tfrac{1}{2}(-x-1)."